OpenStudy (anonymous):

PLEASE! IM STUCK... Evaluate log6 14.

jimthompson5910 (jim_thompson5910):

Hint: Use the change of base formula to get \[\Large \log_{6}(14) = \frac{\log(14)}{\log(6)}\]